OpenStudy (anonymous):

(3)/(x+4)=3 + 4(x-2)

OpenStudy (anonymous):

carry out distributive property on that right side first... can you simplify the right side?

OpenStudy (anonymous):

oops i typed the problem wrong. (3)/(x+4)=3+4/(x-2)

OpenStudy (anonymous):

lol.... ok then... what we need to do here is to clear the fractions by multiplying BOTH SIDES by (x+4)(x-2)...

OpenStudy (anonymous):

\(\large \frac{3}{x+4}=3+\frac{4}{x-2} \) \(\large \color {red} {(x+4)(x-2)}[\frac{3}{x+4}]=\color {red} {(x+4)(x-2)}[3+\frac{4}{x-2}] \) can you simplify the left/right sides?

OpenStudy (anonymous):

i dont understand. can you show me how it looks further

OpenStudy (anonymous):

ok.... let's work with that left side first.... notice the (x+4) will cancell: \(\large \cancel{(x+4)}(x-2)[\frac{3}{\cancel{x+4}}]=3(x-2) \) you follow? this is only the left side...

OpenStudy (anonymous):

i follow!

OpenStudy (anonymous):

great... :) now let's work on the right side...

OpenStudy (anonymous):

the right side is: \(\large (x+4)(x-2)[3+\frac{4}{x-2}] \) from here carry out distributive property to get: \(\large 3(x+4)(x-2)+(x+4)(x-2)\frac{4}{x-2} \) notice the (x-2) will cancel: \(\large 3(x+4)(x-2)+(x+4)\cancel{(x-2)}\frac{4}{\cancel{x-2}} \) can u simplify this?

OpenStudy (anonymous):

is this ok so far?

OpenStudy (anonymous):

i dont know how to simplify that

OpenStudy (anonymous):

i'll help... but do you follow the algebra up to that point?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

ok.... from that last step, \(\large 3(x+4)(x-2)+(x+4)\cancel{(x-2)}\frac{4}{\cancel{x-2}} \) = \(\large 3(x+4)(x-2)+4(x+4) \) = \(\large 3(x^2+2x-8)+4x+16 \) = \(\large 3x^2+6x-24+4x+16 \) = \(\large 3x^2+10x-8 \) remember, this is only the right side... you understand up to this point?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok.... now we cleared the fractions to come up with an equivalent equation to the original. we have the left side = the right side: \(\large 3(x-2)=3x^2+10x-8 \) can you solve from here?

OpenStudy (anonymous):

this is a quadratic equation so solve this like any quadratic by completing the square, factoring, quadratic formula....

OpenStudy (anonymous):

yea thanks!